# Localized Distributional Robustness in Submodular Multi-Task Subset Selection

### v2

@@ 5 Application to Online Submodular Optimization
A natural application of the proposed scheme arises within the context of online submodular optimization, in the presence of time-varying objective functions. In this setting, we aim to produce solutions to a sequence of problems
$\begin{split}&\max_{S_{t}}f_{t}(S_{t})\\ &\text{s.t. }S_{t}\subseteq N,\lvert S_{t}\rvert\leq K,\;\forall t\in\mathbb{N},\end{split}$ (64)
@@
It is easy to approach this problem with the standard methods that we possess. One could treat the problem at each time step $t$ as a separate, standalone problem, and use any variant of the Greedy methods to produce a solution $S_{t},$ entirely decoupled from the problems arising at different time steps.
@@
However, in practical settings, it may be desirable to forgo reconstructing a new solution independent of the previous solutions at each time step, especially if the selection of additional distinct elements incurs extra cost. In this case, it would be desirable to be able to construct a single solution that would perform satisfactorily over multiple time steps, if possible. One strategy to achieve this may be to set an observation window of $t_{w}$ time steps, so that over $t_{w}$ time steps we observe the objective functions played by the system, and only after $t_{w}$ time steps will we play our solution suitable for the $t_{w}$ observations. Ideally, if there is some structure or notion of continuity within the variation of the objective functions $f_{t},$ employing such a strategy might be worthwhile, as it would use up to $t_{w}$ times as few distinct elements while still achieving satisfactory results. In some sense, this would correspond to the inverse of the fairness considerations, where one customarily wants to diversify their selections. Rather, here, the aim is to limit diversification and conserve the same selection of elements over many time steps.
With this idea in mind, one has to consider what sort of strategy to use to leverage the $t_{w}$ observed objective functions $\{f_{1},\ldots,f_{t_{w}}\}$ over a given observation window. Drawing inspiration from prevalent methods that are in use in other optimization areas, e.g., momentum-based stochastic gradient descent methods in continuous optimization [47, 48] and the TD$(\lambda)$ method in reinforcement learning [49], we initially propose the following scheme: For a $\gamma\in[0,1],$
$\begin{split}\max_{S}\gamma^{t_{w}}f_{1}(S)+(1-\gamma)\sum_{t=1}^{t_{w}}\gamma^{(t_{w}-t)}f_{t}(S)\end{split}$ (65)
@@
What this formulation aims to achieve is to geometrically weigh each of the $t_{w}$ observed objective functions, in a manner that may be tuned by the parameter $\gamma$ to determine how much importance will be assigned to older observations with respect to the newer observations. Note that a value of $\gamma=1$ assigns all the weight to the objective function that was observed first, $f_{1}$, whereas $\gamma=0$ assigns, with the convention that $0^{0}=1,$ all the weight to the most recently observed objective function $f_{t_{w}}.$ The values $\gamma\in(0,1)$ provide us with the whole spectrum of how much importance to assign to the past observations.
@@
The important remark to be made in terms of this formulation and the relative entropy regularization is that the geometric weighing of the objective functions over time effectively gives us a reference distribution $\Gamma$, which we can instead use as the center of a neighborhood of robustness, resulting in the formulation
@@
$\begin{split}&\max_{S\subseteq N}\min_{P\in\Delta_{t_{w}}}\sum_{t=1}^{t_{w}}P_{t}f_{t}(S)+\lambda D_{\text{KL}}(P_{t}\;\|\;\Gamma)\\ &\text{s.t. }\lvert S\rvert\leq K,\end{split}$ (66)
@@
where $\Gamma=(\gamma^{t_{w}},(1-\gamma)\gamma^{(t_{w}-2)},\ldots,(1-\gamma)\gamma,1-\gamma)$. This formulation aims to then carry the concept of robustness into the online setting, where we attempt distributional robustness over time steps of the problem.
@@ 6 Experimental Results
Fig. 2: The average performances over fifteen runs of the three algorithms focused on optimizing the reference distribution, the global worst-case task, and the local worst-case distribution as guided by the reference distribution, evaluated on the four criteria of reference distribution performance, worst-case task performance, local worst-case distribution performance and the wall-clock time elapsed taken by the algorithm in the construction of the solution in the satellite selection task of Section 6, Subsection 6.1. Local represents the relative entropy-regularized Stochastic Greedy algorithm solving our novel formulation, aiming for local worst-case distributional robustness in the neighborhood of the reference distribution. Saturate (Global) represents the Submodular Saturation Algorithm proposed in [19], aiming for global worst-case task robustness. Reference represents the Stochastic Greedy algorithm being used to directly optimize the utility of the reference distribution. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with window size 6.
@@
In this section, we will use an application involving a constellation of low Earth orbit (LEO) satellites to test the performance of our three novelly proposed algorithms, namely, the relative entropy-regularized Stochastic Greedy algorithm detailed in Section 4, its application to the online submodular optimization setting detailed in Section 5, and the Saturate with Preference algorithm detailed in Section 3. For the results, we simulate a Walker-Delta constellation parameterized by $i:T/P/f$, where $i$ is the orbit inclination, $T$ is the number of satellites in the constellation, $P$ is the total number of orbital planes of the constellation, and $f$ is the phase difference in between the orbital planes in pattern units[39, 50]. The semi-major axis length of all of the orbits is $8378.1$ kilometers, and the satellites remain Earth-pointing with a field-of-view angle of $\pi/3$ radians. For the testing of all algorithms, we instantiate a constellation with parameters $175^{\circ}:240/12/$ 1.
@@
In the testing of the relative entropy-regularized Stochastic Greedy and the Saturate with Preference algorithms, we instantiate six distinct, time-varying tasks that yield six objective functions $f^{1}_{t},\ldots,f^{6}_{t}$. The objective functions $f^{1}_{t},\ldots,f^{5}_{t}$ relate to the performance of the algorithms on atmospheric sensing tasks. Each of these tasks involves taking atmospheric readings at a set of five randomly located points on Earth. The atmospheric conditions at these points are described by the Lorenz 63 model[51]. The dynamics at these points are parameterized by values that make the system chaotic. An illustration of one instance of these points is provided in Figure 3, where points labeled $1$ belong to atmospheric reading task $1$ whose performance is indicated by the value of $f^{1}_{t}$, and so on. In particular, the utility values $f^{1}_{t}(S),\ldots,f^{5}_{t}(S)$ of these tasks for a given subset of selected satellites $S$ are proportional to the additive inverse of the mean-squared error (MSE) achieved by the selection $S$ of satellites. The MSE is calculated by estimating the state of each atmospheric sensing point using an unscented Kalman filter [52], and comparing it with the actual values of the atmospheric state at the points of interest. This estimation is highly dependent on whether an atmospheric point is within the field of view of a satellite. Note also that the additive inverse of the MSE is known to be a weak-submodular function[53, 54]. Specifically, for $i\in\{1,\ldots 5\}$, $f^{i}_{t}$ has form
$f^{i}_{t}(S_{t})=\frac{1}{Z_{t}^{i}}\operatorname{Tr}\left(\mathbf{P}_{t|t-1}-\mathbf{F}_{S_{t}}^{-1}\right),$ (67)
@@
where $\mathbf{P}_{t|t-1}$ indicates the prediction error covariance matrix at time $t$, $\mathbf{F}_{S_{t}}^{-1}$ indicates the inverse of the Fisher information matrix for the solution $S_{t}$ constructed at time $t$[54], and $Z_{t}^{i}$ is the normalizing coefficient of objective $i$ at time $t$.
@@
Fig. 3: A selection of atmospheric points of interest for the five atmospheric reading tasks, $f^{1},\ldots,f^{5}$ in one run of the simulation. Each task instantiates five points of interest, for a total of twenty-five points. The labels near the points indicate which atmospheric task a point belongs to.
@@
The sixth and final task, whose performance is related by the objective function $f^{6}_{t}$, involves ground coverage of the Earth. The utility value $f^{6}_{t}(S)$ of this task is proportional to the Earth coverage achieved by the selected subset $S$ of the satellites in the constellation. The area of coverage is determined based on a grid of the Earth’s surface consisting of cells of width and height $2^{\circ}$. The area of each cell is calculated assuming a spherical Earth. If the center point of a cell on the grid is within the field-of-view of any satellite in the selection, that grid is taken to be covered by the selection. Specifically, $f^{6}_{t}$ has form
$f^{6}_{t}(S_{t})=\dfrac{1}{Z^{6}_{t}}\dfrac{\sum_{c=1}^{C}\mathbb{1}(\text{cell }c\text{ is in view of a satellite in }S_{t})}{C},$ (68)
where $C$ is the total number of grid-cells, and $Z^{6}_{t}$ is the normalizing coefficient of objective $6$ at time $t$.
@@
As indicated above, the utility values of all six tasks $f^{1},\ldots,f^{6}$ are normalized to the range of $[0,1]$, by dividing the utility $f^{i}(S)$ of a selection $S$ on task $f^{i}$ at any time step by $f^{i}(N)$, the maximum utility achievable by selecting the entire ground set at that time step. This ensures that the importance score of an individual task $f^{i}$ is not influenced by the potentially arbitrary value of its utility, and is solely dictated by the weight $Q_{i}$ assigned to it.
@@
We randomly sample a reference distribution $Q$ uniformly from $\Delta_{n}$ to assign an importance score $Q_{i}$ to each task $f^{i}$. Following, we simulate the atmospheric states of the points of interest and the trajectory of the Walker-Delta constellation, for 25 time steps, each corresponding to a time interval of 60 seconds, for 15 runs.
6.1 Relative Entropy-Regularized Stochastic Greedy for Satellite Selection
@@
For the assessment of the relative entropy-regularized Stochastic Greedy algorithm, we compare the performance of the solutions produced by three algorithms, in terms of four distinct performance criteria. For each of the algorithms, we use a cardinality constraint of $K=10$, a sampling set size of $|R|=24$, and a regularization parameter of $\lambda=0.1$, where applicable. The three algorithms compared are as follows:
@@
  • Algorithm 1 - Local: The relative-entropy regularized Stochastic Greedy algorithm, that is, the Stochastic Greedy algorithm applied to the maximization of our novel objective function $g(S)=-\lambda\log(\sum_{i=1}^{n}Q_{i}\exp{(-f^{i}(S)/\lambda)}),$ of Problem (25).
@@
  • Algorithm 2 - Saturate (Global): The Submodular Saturation Algorithm, which aims to optimize the worst-case scenario in Problem (2).
@@
  • Algorithm 3 - Reference: The Stochastic Greedy algorithm, applied to the maximization of the objective function $f(S)=\sum_{i=1}^{n}Q_{i}f^{i}(S)$ of Problem (4), focused on directly optimizing with respect to the reference distribution with no consideration of robustness.
@@
  • Criterion 1: The utility scores with respect to the objective $F_{1}(S)=\sum_{i=1}^{n}Q_{i}f_{i}(S)$, which represents the performance of the algorithms when weighed directly by the reference distribution itself.
  • Criterion 2: The utility scores with respect to the objective $F_{2}(S)=\min_{i}f_{i}(S)$, which represents the global worst-case single task performance.
  • Criterion 3: The utility scores with respect to the objective $F_{3}(S)=\sum_{i=1}^{n}P^{*}_{i}f_{i}(S)$, where $P^{*}=\arg\!\min_{P}\sum^{n}_{i=1}P_{i}f^{i}(S)+\lambda D_{\text{KL}}(P\;\|\;Q)$, hence representing the local worst-case scenario performance localized to a relative entropy neighborhood of the reference distribution. This criterion may be considered as the benchmark of our novel formulation.
  • Criterion 4: The runtime of the algorithm.
@@
Several observations may be made looking at these results. Firstly, the designation of Saturate as “too pessimistic” is indeed justified by the results, as seen from Figure 2 (a), where it fails to perform on the reference distribution as it is too focused on worst-case performance. However, it does indeed dominate in worst-case single-task performance over the other two algorithms, as evidenced by Figure 2 (b).
@@
Concerning the performance of Algorithm 1 - Local solving our novel formulation, we see from Figure 2 (a) that it has comparable performance on the reference distribution with Algorithm 3 - Reference, which directly optimizes the reference distribution. The usefulness of the Local approach is reflected clearly in Figure 2 (c), where we observe it fulfilling the local distributional robustness that our novel formulation aimed at achieving. Algorithm 1 - Local greatly outperforms Algorithm 3 - Reference and marginally outperforms Saturate on the worst-case scenario within a neighborhood around the reference distribution. However, this marginal surpassing is complemented by the much smaller runtime and computational complexity of the Local approach.
With regard to the wall-clock time taken by the algorithms, as expected, Algorithms 1 and 3 perform identically, as they are essentially the same Stochastic Greedy algorithm. Saturate performs much more poorly, consistently taking many times as much time at each time step. This is explained by the fact that due to its line search-based approach, Saturate uses several runs of the Stochastic Greedy procedure in each of its iterations.
We may conclude from these results that Algorithm 1 - Local, solving our proposed novel formulation, succeeds in constructing a solution that is comparable to Algorithm 3 - Reference in optimizing the performance of the reference distribution, while also achieving local distributional robustness within the neighborhood of the reference distribution. Although Saturate also manages to produce a solution that is more robust against worst-case tasks and achieves decent local distributional robustness, it is much more computationally expensive and has a much longer runtime.

### v3

@@ 5 Application to Online Submodular Optimization
A natural application of the proposed scheme arises within the context of online submodular optimization, in the presence of time-varying objective functions. In this setting, we aim to produce a sequence of solutions to a sequence of problems
$\begin{gathered}\max_{\mathcal{S}_{t}\in\mathcal{N}}f_{t}(\mathcal{S}_{t})\\ \text{s.t.}\;\lvert\mathcal{S}_{t}\rvert\leq K,\;\forall t\in\mathbb{N},\end{gathered}$ (64)
@@
It is easy to approach this problem with the standard methods that we possess. We could treat the problem at each time step $t$ as a separate, standalone problem, and use any variant of the Greedy methods to produce a solution $\mathcal{S}_{t},$ entirely decoupled from the problems arising at different time steps.
@@
However, in practical settings, it may not be desirable to reconstruct a new solution independent of the previous solutions at each time step, especially if the selection of additional distinct elements incurs extra cost. In this case, it would be desirable to construct a single solution that would perform satisfactorily over multiple time steps, if possible. One strategy to achieve this may be to set an observation window of $t_{w}$ time steps, so that over $t_{w}$ time steps we observe the objective functions played by the system, and only after $t_{w}$ time steps will we play our solution suitable for the $t_{w}$ observations. Ideally, if there is some structure or notion of continuity within the variation of the objective functions $f_{t},$ employing such a strategy might be worthwhile, as it would use up to $t_{w}$ times as few distinct elements while still achieving satisfactory results. In some sense, this is like an inverse of fairness considerations, where one customarily wants to diversify their selections. Rather, here, the aim is to limit diversification and conserve the same selection of elements over many time steps.
[…]
$\begin{split}\max_{\mathcal{S}\in\mathcal{N}}\gamma^{t_{w}}f_{1}(\mathcal{S})+(1-\gamma)\sum_{t=1}^{t_{w}}\gamma^{(t_{w}-t)}f_{t}(\mathcal{S})\end{split}$ (65)
@@
What this formulation aims to achieve is to geometrically weigh each of the $t_{w}$ observed objective functions, where $\gamma$ decides the relative importance of older observations with respect to the newer observations. A value of $\gamma=1$ assigns all the weight to $f_{1}$, whereas $\gamma=0$ assigns, with the convention that $0^{0}=1,$ all the weight to the most recently observed objective function $f_{t_{w}}.$ The values $\gamma\in(0,1)$ provide us with the whole spectrum of relative importances.
@@
One important remark is that the geometric weighing of the objective functions over time effectively gives us a reference distribution $\Gamma$, which we can instead use as the center of a neighborhood of robustness, resulting in the formulation
@@
$\begin{gathered}\max_{\mathcal{S}\subseteq\mathcal{N}}\min_{P\in\Delta_{t_{w}}}\sum_{t=1}^{t_{w}}P_{t}f_{t}(\mathcal{S})+\lambda D_{\text{KL}}(P_{t}\;\|\;\Gamma)\\ \text{s.t.}\;\lvert\mathcal{S}\rvert\leq K,\end{gathered}$ (66)
@@
where $\Gamma=((1-\gamma)\gamma^{t_{w}-1},\ldots,(1-\gamma)\gamma,(1-\gamma))$. This formulation aims to then carry the concept of robustness into the online setting, where we attempt distributional robustness over time steps of the problem.
@@ 6 Experimental Results
Fig. 2: The average performances over $15$ runs of the three algorithms focused on optimizing the reference distribution, the global worst-case task, and the local worst-case distribution close to the reference distribution, evaluated on the four criteria of reference distribution performance, worst-case task performance, local worst-case distribution performance and the wall-clock time taken by the algorithm in the construction of the solution in the satellite selection task of Section 6.1. Local represents the relative-entropy-regularized Stochastic Greedy algorithm solving our novel formulation, aiming for local worst-case distributional robustness in the neighborhood of the reference distribution. Saturate (Global) represents SSA proposed in [35], aiming for global worst-case task robustness. Reference represents the Stochastic Greedy algorithm being used to directly optimize the utility of the reference distribution. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with window size 66.
@@
In this section, we use an application involving a constellation of low Earth orbit (LEO) satellites to test the performance of our three novelly proposed algorithms, namely, the relative-entropy-regularized Stochastic Greedy algorithm detailed in Section 4, its application to the online submodular optimization setting detailed in Section 5, and the Saturate with Preference algorithm detailed in Section 3. For the results, we simulate a Walker-Delta constellation parameterized by $i:T/P/f$, where $i$ is the orbit inclination, $T$ is the number of satellites in the constellation, $\Pi$ is the total number of orbital planes of the constellation, and $\phi$ is the phase difference in between the orbital planes in pattern units[25, 56]. The semi-major axis length of all of the orbits is $8378.1$ kilometers, and the satellites remain Earth-pointing with a field-of-view angle of $\pi/3$ radians. For the testing of all algorithms, we instantiate a constellation with parameters $75^{\circ}:240/12/1$.
@@
In the testing of the relative-entropy-regularized Stochastic Greedy and the Saturate with Preference algorithms, we instantiate six distinct, time-varying tasks that yield six objective functions $f_{t,1},\ldots,f_{t,6}$. The objective functions $f_{t,1},\ldots,f_{t,5}$ are related to the performance of the algorithms on atmospheric sensing tasks. Each of these tasks involves taking atmospheric readings at a set of five randomly located points on Earth. The atmospheric conditions at these points are described by the Lorenz 63 model[41]. The dynamics at these points are parameterized by values that make the system chaotic. An illustration of one instance of these points is provided in Figure 3, where points labeled $1$ belong to atmospheric reading task $1$ whose performance is indicated by the value of $f_{t,1}$, and so on. In particular, the utility values $f_{t,1}(\mathcal{S}),\ldots,f_{t,5}(\mathcal{S})$ of these tasks for a subset of selected satellites $\mathcal{S}$ are proportional to the additive inverse of the mean-squared error (MSE) achieved by the selection $\mathcal{S}$ of satellites. The MSE is calculated by estimating the state of each atmospheric sensing point using an unscented Kalman filter [3], and comparing it with the actual values of the atmospheric state at the points of interest. This estimation is highly dependent on whether an atmospheric point is within the field of view of a satellite. Note also that the additive inverse of the MSE is known to be a weak-submodular function[24, 22]. Specifically, for $i\in[5]$, $f_{t,i}$ has form
$f_{t,i}(\mathcal{S}_{t})=\frac{1}{Z_{t,i}}\operatorname{Tr}\left(\mathbf{P}_{t|t-1}-\mathbf{F}_{\mathcal{S}_{t}}^{-1}\right),$ (67)
@@
where $\mathbf{P}_{t|t-1}$ indicates the prediction error covariance matrix at time $t$, $\mathbf{F}_{\mathcal{S}_{t}}^{-1}$ indicates the inverse of the Fisher information matrix for the solution $\mathcal{S}_{t}$ constructed at time tt[22], and $Z_{t,i}$ is the normalizing coefficient of objective $i$ at time $t$.
@@
Fig. 3: A selection of atmospheric points of interest for the five atmospheric reading tasks, $f_{t,1},\ldots,f_{t,5}$ in one run of the simulation. Each task instantiates 55 points of interest, for a total of $25$ points. The labels near the points indicate which atmospheric task a point belongs to.
@@
The sixth and final task, whose performance is given by the objective function $f_{t,6}$, involves ground coverage of the Earth. The utility value $f_{t,6}(\mathcal{S})$ of this task is proportional to the Earth coverage achieved by the selected subset $\mathcal{S}$ of the satellites in the constellation. The area of coverage is determined based on a grid of the Earth’s surface consisting of cells of width and height $2^{\circ}$. The area of each cell is calculated assuming a spherical Earth. If the center point of a cell on the grid is within the field-of-view of any satellite in the selection, that grid is taken to be covered by the selection. Specifically, $f_{t,6}$ has form
[…]
where $C$ is the total number of grid-cells.
@@
The utility values of all six tasks $f_{t,1},\ldots f_{t,6}$ are normalized to the range of $[0,1]$, by dividing the utility $f_{ti}(\mathcal{S})$ of a selection $\mathcal{S}$ at any time step by $t$ $f_{t,i}(\mathcal{N})$, the maximum utility achievable by selecting the entire ground set at that time step. This ensures that the importance score of an individual task is not influenced by the potentially arbitrary value of its utility, and is solely determined by the weight $Q_{i}$ assigned to it.
@@
We randomly sample a reference distribution $Q$ uniformly from $\Delta_{n}$ to assign an importance score to each task. Following, we simulate the atmospheric states of the points of interest and the trajectory of the Walker-Delta constellation, for $25$ time steps, each corresponding to a time interval of $60$ seconds, for $15$ runs.
6.1 Relative-Entropy-Regularized Stochastic Greedy for Satellite Selection
@@
For the assessment of the relative-entropy-regularized Stochastic Greedy algorithm, we compare the performance of the solutions produced by three algorithms, in terms of four distinct performance criteria. For each of the algorithms, we use a cardinality bound $K=10$, a sampling set size $\lvert\mathcal{R}\rvert=24$, and a regularization parameter $\lambda=0.1$, where applicable. The three algorithms compared are as follows:
@@
  • Algorithm 1 - Local: The relative-entropy-regularized Stochastic Greedy algorithm, that is, the Stochastic Greedy algorithm applied to the maximization of our novel objective function $G(\mathcal{S})=-\lambda\log(\sum_{i=1}^{n}Q_{i}\exp{(-f_{i}(\mathcal{S})/\lambda)})$ in (25).
@@
  • Algorithm 2 - Saturate (Global): The Submodular Saturation Algorithm, which aims to optimize the worst-case scenario of (2).
@@
  • Algorithm 3 - Reference: The Stochastic Greedy algorithm, applied to the maximization of the objective function $f(\mathcal{S})=\sum_{i=1}^{n}Q_{i}f_{i}(\mathcal{S})$ of (4), focused on directly optimizing with respect to the reference distribution with no consideration of local robustness.
@@
  • Criterion 1: The utility scores with respect to the objective $F_{1}(\mathcal{S})=\sum_{i=1}^{n}Q_{i}f_{i}(\mathcal{S})$, which represents the performance of the algorithms when weighed directly by the reference distribution $Q$.
  • Criterion 2: The utility scores with respect to the objective $F_{2}(\mathcal{S})=\min_{i\in[n]}f_{i}(\mathcal{S})$, which represents the global worst-case single-task performance.
  • Criterion 3: The utility scores with respect to the objective $F_{3}(\mathcal{S})=\sum_{i=1}^{n}P^{*}_{i}f_{i}(\mathcal{S})$, where $P^{*}=\arg\!\min_{P\in\Delta_{n}}\sum^{n}_{i=1}P_{i}f_{i}(\mathcal{S})+\lambda D_{\text{KL}}(P\;\|\;Q)$, hence representing the local worst-case scenario performance localized to a relative-entropy neighborhood of the reference distribution. This criterion may be considered as the benchmark of our novel formulation.
  • Criterion 4: The wall-clock runtime of the algorithm.
@@
Several observations follow from these results. Firstly, the designation of SSA as “too pessimistic” is indeed justified by the results, as seen from Figure 2 (a), where it fails to perform on the reference distribution as it is too focused on worst-case performance. However, it does indeed dominate in worst-case single-task performance over the other two algorithms, as evidenced by Figure 2 (b).
@@
Concerning the performance of Algorithm 1 - Local solving our formulation, we see from Figure 2 (a) that it has comparable performance on the reference distribution with Algorithm 3 - Reference, which directly optimizes the reference distribution. The usefulness of the Local approach is reflected clearly in Figure 2 (c), where we observe it fulfilling the local distributional robustness that our formulation aimed at achieving. Algorithm 1 - Local greatly outperforms Algorithm 3 - Reference and marginally outperforms SSA on the worst-case scenario within a neighborhood around the reference distribution. However, this marginal surpassing is complemented by the much smaller runtime and computational complexity of the Local approach.
With regard to the wall-clock time taken by the algorithms, as expected, Algorithms 1 and 3 perform identically, as they are instances of the same Stochastic Greedy algorithm. SSA performs much more poorly, consistently taking many times as much time at each time step. This is explained by the fact that due to its line search-based approach, SSA uses several runs of the Stochastic Greedy procedure in each of its iterations.
We may conclude from these results that Algorithm 1 - Local, solving our proposed novel formulation, succeeds in constructing a solution that is comparable to Algorithm 3 - Reference in optimizing the performance of the reference distribution, while also achieving local distributional robustness within the neighborhood of the reference distribution. Although SSA also manages to produce a solution that is more robust against worst-case tasks and achieves decent local distributional robustness, it is much more computationally expensive and has a much longer runtime.